OpenStudy (anonymous):

The sum of 5 times a number and 5 is twice the number. Find the number.

OpenStudy (anonymous):

Can someone please tell me how to write this problem out

OpenStudy (mary.rojas):

5x+5=2x

OpenStudy (mary.rojas):

solve x

OpenStudy (anonymous):

hmm

OpenStudy (mary.rojas):

... I think that is how. But have some one check

OpenStudy (anonymous):

5x+5=2x

OpenStudy (anonymous):

-5/3????????

OpenStudy (mary.rojas):

plug it in and see if it makes sense

OpenStudy (anonymous):

-5/3 = -1.66

OpenStudy (mary.rojas):

Yes it makes sense. do you see why?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hmm please explain

OpenStudy (mary.rojas):



OpenStudy (mary.rojas):

so they equal each other when x=-5/3

OpenStudy (anonymous):

oh ok I get it!!!! thank you

OpenStudy (mary.rojas):

your welcome.